# Transforming Code Repositories into MCP Services

### v3

@@ 4.3 Comparison with Human and GPT-4 Baselines
Figure 4: Repo success rates with and without Run‑Review‑Fix (left) and remaining errors across RRF rounds (right) for five representative scientific MCP repositories.
@@
In the human configuration, ten graduate students with at least three years of Python experience and basic familiarity with MCP concepts implement wrappers from scratch. Before the study, all participants read the MCP specification and a commented example repository. Each participant is randomly assigned several repositories, and the measured time is from the moment they start reading the repository to the point where the minimal test passes.
@@
In the GPT-4 template configuration, a unified system prompt instructs a single GPT-4 model to generate a complete MCP service implementation based on the repository README and several key source files selected by static analysis. We use the same base model and decoding configuration as in Code2MCP, namely gpt-4o-2024-05-13 with temperature set to 1.0, to ensure a fair comparison. The model is allowed up to three dialogue turns: the first turn generates an initial implementation, and up to two additional turns may revise the code based on execution errors returned by a test runner. No additional agents or planning mechanisms are used, and the full system prompt is provided in Appendix B.1.
Figure 5: Average task success rate (left) and average completion time (right) across the ten scientific domains for the three configurations: Human experts, GPT-4 with basic scaffolding, and Code2MCP.
@@
In the Code2MCP configuration, the default multi-agent pipeline with the Run-Review-Fix loop is used. For each repository, we record the final success status and the completion time, defined as the end-to-end wall-clock time starting from the beginning of environment setup and ending when the unified MCP test first passes or the run is terminated; see Appendix B.2 for details. Figure 5 summarizes the average task success rates and completion times of the three configurations across the ten scientific domains. For each configuration and domain, task success rate is defined as the fraction of successful attempts over all attempts on representative repositories in that domain. Completion time is the average end-to-end wall-clock time per attempt, measured from the beginning of environment setup until the minimal MCP test first passes or the run is terminated.
Figure 6: Task success rates of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups. The x-axis lists representative repositories in each group, labeled by domain abbreviation and index (e.g., Math-1).
The overall trend is as follows: Code2MCP achieves task success rates close to human implementations in most domains, and consistently higher than the GPT-4 template configuration. In terms of completion time, human implementations are typically on the order of hours, while Code2MCP completes in tens of minutes; the GPT-4 template configuration is the fastest but has lower success rates.
When repositories are grouped by dependency complexity and project type into low-dependency libraries, medium-dependency scientific libraries, high-dependency libraries, and script-heavy projects, Figures 6 and 7 further break down the same comparison on representative repositories within each group. The bucketing rules are detailed in Appendix B.2.
Figure 7: Completion time of Human experts, GPT-4 with a template, and Code2MCP on representative repositories in the four repository groups.
4.4 Integration with Existing Tool Systems
@@
RepoMaster + Code2MCP. In the RepoMaster setting, code-understanding and tool-usage tasks are run in two configurations: using RepoMaster alone to operate at the source-code level, and using RepoMaster together with Code2MCP, where RepoMaster first selects relevant repositories and then invokes the corresponding MCP tools. Figure 8 shows that the configuration with Code2MCP achieves higher task success rates and requires fewer interaction steps (see Appendix B.2 for metric definitions), suggesting that delegating part of the work to standardized MCP tools improves efficiency without modifying the upstream agent policy.
@@
OpenAgents + Code2MCP. In the OpenAgents setting, cross-domain tasks are evaluated under two configurations: using OpenAgents alone, and using OpenAgents with MCP tools automatically generated by Code2MCP added to the tool pool. As shown in Figure 9, we report the task success rate and coverage, where coverage and success rate are defined in Appendix B.2. Adding Code2MCP tools increases both metrics across all task groups, indicating that a richer MCP tool pool can be effectively exploited by existing planning and retrieval strategies.
@@
Overall, RepoMaster and OpenAgents handle tool discovery and code retrieval on the consumer side, while Code2MCP supplies additional MCP services on the supply side. Under the same front-end selection and planning strategies, integrating Code2MCP enriches the tool pool and yields measurable improvements in end-to-end task performance.
Figure 8: RepoMaster with and without Code2MCP tools: task success rate (left) and average interaction steps (right) across task groups.
4.5 Case Studies: Protein, Math, and Computational Fluid Dynamics
@@
This subsection highlights three representative repositories from the 50-repository evaluation set: the biomedical protein modeling library ESM, the symbolic mathematics library SymPy, and the CFD framework Foam-Agent built on OpenFOAM, illustrating the kinds of MCP tools Code2MCP generates and how agents use them in practice.
Figure 9: OpenAgents with and without Code2MCP tools: task success rate (left) and coverage (right) across general and scientific task groups.
@@
Bridging protein science with AI agents. In protein science, models such as AlphaFold (Jumper et al., 2021) greatly improve structure prediction, but using them in everyday research typically requires substantial scripting and environment setup. ESM models complement AlphaFold with efficient sequence modeling and zero-shot variant effect prediction, yet traditionally still require multi-step Python pipelines for computing physicochemical properties, predicting structures, and analyzing mutations. Code2MCP converts ESM into an MCP service by exposing tools such as Analyze_sequence, Predict_structure, and Predict_variant_effect. We provide a qualitative example of using ESM-based MCP tools for protein analysis in Appendix A.5.
Figure 10: An AI agent orchestrates a CFD simulation pipeline by invoking Generate_mesh, Generate_simulate, and Visualize_velocity functions to guide a user through the entire process, from initial mesh generation and simulation setup to the final velocity visualization.
@@
Enhancing mathematical reliability in AI agents. For mathematical computing, large language models are prone to errors in symbolic derivations and exact calculations, whereas libraries such as SymPy already provide reliable implementations. Code2MCP organizes common capabilities from SymPy into a mathematical MCP service with tools for limits and integration, matrix operations, symbolic simplification, and transforms such as Fourier and Laplace. When asked to compute an integral or the volume of a solid of revolution, an agent simply invokes the corresponding tool and returns an exact answer, relying on SymPy as a backend instead of ad-hoc prompt-based reasoning.
@@
Automating CFD simulation for AI agents. CFD workflows typically involve mesh generation, solver configuration, execution, and post-processing, traditionally coordinated by engineers through shell scripts and configuration files. Foam-Agent provides a higher-level interface to OpenFOAM, and Code2MCP further turns it into an MCP service with tools such as Generate_mesh, Generate_simulate, and Visualize_velocity. As illustrated in Figure 10, a user describes the target geometry and physical parameters in natural language, and the agent calls these tools in sequence to construct the mesh, run the simulation, and visualize the velocity field.
5 Conclusion
@@
Addressing the key challenge of insufficient tool supply in the MCP ecosystem, this paper introduces Code2MCP, a framework that automatically converts GitHub repositories into functional MCP services. The framework employs a multi-agent workflow for code comprehension, environment reconstruction, and tool abstraction, augmented by a “Run-Review-Fix” self-correction loop that improves end-to-end reliability across diverse scientific libraries. Our evaluation on 50 repositories shows that Code2MCP can robustly expose high-value APIs as MCP tools while also revealing clear failure modes and boundary conditions. This work presents an automated pathway for connecting existing codebases to the agent tool ecosystem, and we leave systematically hardening security and building standardized large-scale MCP benchmarks from these converted services as important directions for future work.
@@ Appendix A Appendix
A.1 Use of Large Language Models
@@
A.2 Output Directory Structure
Figure 11 shows the final directory layout generated by Code2MCP, where the original repository and the synthesized MCP artifacts are organized under a unified workspace.
Figure 11: The complete output directory structure generated by the Code2MCP framework. The top-level workspace contains the original repository alongside all generated artifacts within the mcp_output directory.
A.3 Detailed Conversion Pipeline and Algorithm
This section outlines the roles of the specialized agents within the Code2MCP framework and their coordination as shown in Algorithm 1, where Download and Environment initialize the workspace, Analysis and Generation construct the MCP service, Run and Review form the Run-Review-Fix loop, and Finalize produces the final deliverables. The following are placeholders for their system prompts.
@@ A.3 Detailed Conversion Pipeline and Algorithm
A.4 Agent Roles and System Prompts
This section outlines the roles of the specialized agents within the Code2MCP framework. The following are placeholders for their system prompts.
Environment System Prompt - Prefer Conda; use venv only if Conda is unavailable or clearly unsuitable. - Detect dependency sources by priority: environment.yml > requirements.txt > pyproject.toml > setup files; never guess hidden dependencies. - Pin versions when explicit; otherwise install the minimal viable set. Prefer CPU wheels unless GPU is explicitly required. - Normalize cross-platform behavior; avoid absolute paths; use relative POSIX-like paths; ensure UTF-8 locale. - Smoke test: print Python version and platform; import fastmcp; attempt to import the project’s top-level package or a primary CLI; exit code 0 indicates pass. - On failure, capture exact command, exit code, last 80 lines of stderr, and timing; propose exactly one minimal remedy (e.g., switch to venv, install a single missing package, try one version pin, extend timeout once). - Apply at most one fallback; never change repository code; do not write outside the workspace; do not weaken security (e.g., no SSL bypass). - Cache wheels where possible; avoid global pollution; record reproducible commands and resolved versions. - Default to offline validation; if network is strictly required, justify briefly and bound the scope. - Emit a compact environment report (platform, Python, manager, explicit deps, resolved pins, pass/fail).
Review & Auto-Fix System Prompt - Triage failures: import/env, type/contract, path/I-O, dependency/version, timeout/perf, platform. - Choose minimal direct fix vs. regeneration; provide a one-line rationale and confidence (low/med/high). Prefer adapter-boundary mitigations (lazy import, existence checks, parameter coercion). - Apply strict complete-file replacement for the single target file; return pure code only; do not alter unrelated sections. - Preserve interface contracts and standardized error shapes; add narrow guards instead of broad catches. - Enforce cross-platform path handling and deterministic behavior; do not introduce external network calls or new side effects. - Optionally add a tiny internal sanity check if it prevents recurrence without bloat. - Bound attempts (<= B). If still failing, emit a compact escalation note: failing step, last command, stderr tail, and the next best single remediation.
A.5 Additional Protein Case Study: ESM
Figure 12: An AI agent processes a user’s query containing a protein sequence and mutation. By invoking Analyze_sequence, Predict_structure, and Predict_variant_effect functions from the generated ESM MCP server, it returns key physicochemical properties, a predicted 3D structure, and an analysis of the mutation’s effects.
Appendix B Details of Evaluation
B.1 GPT-4 Template Baseline Prompt
B.2 Evaluation Metrics and Bucketing Rules
This section defines the evaluation metrics and repository bucketing rules used in the main experiments.
Interaction steps. In the RepoMaster experiments, an interaction step is defined as a single high-level agent action that changes the external state of the system. Concretely, we count as one step any of the following: (1) invoking a tool (including Code2MCP-generated MCP tools and built-in RepoMaster tools), (2) editing source code files, or (3) executing the target program or tests. The total number of interaction steps for a task is the length of the resulting action sequence.
Task Set Description. In the integration experiments with RepoMaster and OpenAgents, we use a fixed set of code-understanding and tool-usage tasks that are primarily derived from public benchmarks and representative scenarios in prior work, with a small number of additional scientific-computing cases to ensure coverage of both general and scientific settings. At this stage, this task set remains a relatively small, research-oriented setup rather than a fully standardized large-scale benchmark, and constructing a more systematic MCP task benchmark is left as future work.
Task success rate and coverage. In the OpenAgents experiments, a task is considered successful if the agent produces a final answer that passes automatic correctness checks. Task success rate is the fraction of tasks in a group that are successful. Coverage is the fraction of tasks in a group for which the agent produces any final answer that can be checked by the automatic evaluator, regardless of whether the answer is ultimately judged correct.
  • Low-dependency libraries: at most 10 Python dependencies inferred from requirements.txt, environment.yml, or pyproject.toml, and primarily library-style public APIs.
@@ Appendix C Per-repository Results
Table 3: Repository identifiers, domains, and final outcomes.
@@
B1 Biomedical facebookresearch/esm success
@@
B4 Biomedical pysam-developers/pysam success
B5 Biomedical deepchem/deepchem success
@@
P2 Psychology pymc-devs/pymc success
P3 Psychology sahahn/BPt success
@@
P5 Psychology neuropsychology/NeuroKit success
M1 Math sympy/sympy success
@@
M3 Math fredrik-johansson/mpmath success
M4 Math cvxpy/cvxpy success
M5 Math sagemath/sage success
Ea1 Earth Science obspy/obspy success
@@
Ea4 Earth Science mapbox/rasterio success
Ea5 Earth Science geopandas/geopandas success
A1 Astronomy astropy/astropy success
@@
A3 Astronomy lightkurve/lightkurve success
A4 Astronomy astroML/astroML success
A5 Astronomy astropy/astroquery success
C1 Chemistry rdkit/rdkit success
@@
C3 Chemistry bjodah/chempy success
@@
C5 Chemistry cclib/cclib success
Ph1 Physics csml-rpi/Foam-Agent success
@@
Ph3 Physics pydy/pydy fail
Ph4 Physics PyAbel/PyAbel success
@@
S1 Social Science networkx/networkx success
S2 Social Science igraph/python-igraph success
@@
S5 Social Science datamade/dedupe success
L1 Linguistics nltk/nltk success
L2 Linguistics explosion/spaCy success
L3 Linguistics stanfordnlp/stanza success
@@
L5 Linguistics flairNLP/flair success
E1 Econometrics statsmodels/statsmodels success
E2 Econometrics alkaline-ml/pmdarima success
@@
E5 Econometrics pydata/xarray success
Table 4: Per-repository structural statistics (size, dependencies, and tests).
@@
Table 5: Per-repository environment and MCP conversion statistics.
@@
Appendix D Example MCP Tool Implementations Generated by Code2MCP
D.1 Tools Generated from the ESM Repository
MCP Tool for Protein Sequence Analysis @mcp.tool(name="analyze_sequence", description="Analyze protein sequence features.") def analyze_sequence(sequence: str): """Analyzes physicochemical properties of a protein sequence.""" try: import re try: from esm import analysis except Exception: try: from .esm import analysis except Exception: import types analysis = types.SimpleNamespace( calculate_molecular_weight=lambda s: float(len(s)) * 110.0, calculate_isoelectric_point=lambda s: 7.0, calculate_instability_index=lambda s: 40.0 ) aa_set = set("ACDEFGHIKLMNPQRSTVWY") seq = re.sub(r"[^A-Za-z]", "", sequence or "").upper() seq = "".join([c for c in seq if c in aa_set]) length = len(seq) composition = {$a$: seq.count($a$) for $a$ in aa_set if seq.count($a$) > 0} molecular_weight = analysis.calculate_molecular_weight(seq) isoelectric_point = analysis.calculate_isoelectric_point(seq) instability_index = analysis.calculate_instability_index(seq) properties = { "length": length, "composition": composition, "molecular_weight": molecular_weight, "isoelectric_point": isoelectric_point, "instability_index": instability_index, "sequence": seq, } return {"success": True, "result": properties, "error": None} except Exception as e: return {"success": False, "result": None, "error": f"Error during sequence analysis: {str(e)}"}
[…]
D.1.1 Tools Generated from the SymPy Repository
D.1.2 Tools Generated from the Foam-Agent Repository

### v4

@@ 4.3. Comparison with Human and GPT-4 Baselines
In the human configuration, ten graduate students with at least three years of Python experience and basic familiarity with MCP concepts implement wrappers from scratch. Before the study, all participants read the MCP specification and a commented example repository. Each participant is randomly assigned several repositories. For consistency, the same success criterion and the same validation procedure are applied. Completion time is measured from the start of wrapper implementation to the first time the unified validation passes, or to termination under the same budgets.
@@
In the GPT-4 template configuration, a unified system prompt instructs a single model to generate a complete MCP service implementation based on the repository README and several key source files selected by static analysis. The model performs end-to-end interface design, code generation, execution, and iterative debugging based on execution tracebacks, constituting a single-agent baseline. The baseline follows the same validation procedure and the same success criterion. Generation and revision are performed iteratively based on execution errors returned by the test runner, and the process stops when the unified validation passes or when the evaluation budget is exhausted. No additional agents, retrieval modules, or external planning tools are used. The full system prompt is provided in Appendix E.1.
(a) RepoMaster with/without Code2MCP.
(b) OpenAgents with/without Code2MCP.
Figure 6. (a) shows RepoMaster with and without Code2MCP tools, reporting task success rates and average interaction steps across task groups. (b) shows OpenAgents with and without Code2MCP tools, reporting task success rates and coverage across general and scientific task groups.
@@
In the Code2MCP configuration, the default multi-agent pipeline with the Run-Review-Fix loop is used. For each repository, we record the final success status and the completion time, defined as the end-to-end wall-clock time starting from the beginning of environment setup and ending when the unified MCP test first passes or the run is terminated. To make the evaluation protocol fully self-contained.
Figure 7 summarizes the average task success rates and completion times of the three configurations across the ten scientific domains. For each configuration and domain, task success rate is defined as the fraction of successful attempts over all attempts on representative repositories in that domain. Completion time is the average end-to-end wall-clock time per attempt, measured from the beginning of environment setup until the minimal MCP test first passes or the run is terminated. The overall trend is as follows: Code2MCP achieves task success rates close to human implementations in most domains, and higher than the GPT-4 template configuration. In terms of completion time, human implementations are on the order of hours, while Code2MCP completes in tens of minutes; the GPT-4 template configuration is the fastest but has lower success rates. When repositories are grouped by dependency complexity and project type into low-dependency libraries, medium-dependency scientific libraries, high-dependency libraries, and script-heavy projects, Figures 5 further break down the same comparison on representative repositories within each group.
4.4. Integration with Existing Tool Systems
@@
Task Set Description. In the integration experiments with RepoMaster and OpenAgents, we evaluate on a fixed set of code-understanding and tool-usage tasks designed to assess whether automatically generated MCP tools can be discovered, selected, and invoked by existing agent systems under realistic usage patterns. The task categories and evaluation protocol are kept identical across methods, and full task specifications and examples are provided in Appendix D.
@@
RepoMaster + Code2MCP. In the RepoMaster setting, tool-usage and code-understanding tasks are run in two configurations: using RepoMaster alone to operate at the source-code level, and using RepoMaster together with Code2MCP, where RepoMaster first selects relevant repositories and then invokes the corresponding MCP tools. Figure 6(a) shows that the configuration with Code2MCP achieves higher task success rates and requires fewer interaction steps, suggesting that delegating part of the work to standardized MCP tools improves efficiency.
Figure 7. Average task success rate (left) and average completion time (right) across the ten scientific domains for the three configurations: Human experts, GPT-4 with basic scaffolding, and Code2MCP.
OpenAgents + Code2MCP. In the OpenAgents setting, cross-domain tasks are evaluated under two configurations: using OpenAgents alone, and using OpenAgents with MCP tools automatically generated by Code2MCP added to the tool pool. A task is considered successful if the agent produces a final answer that passes automatic correctness checks.
(a) ESM protein analysis workflow.
(b) Foam-Agent CFD workflow.
Figure 8. (a) shows a protein analysis pipeline where an agent invokes Analyze_sequence, Predict_structure, and Predict_variant_effect from the generated ESM MCP server to return physicochemical properties, a predicted 3D structure, and mutation impact analysis. (b) shows a CFD pipeline where the agent calls Generate_mesh, Generate_simulate, and Visualize_velocity to guide the user from mesh generation and solver setup through simulation execution and final velocity-field visualization.
@@
Task success rate is the fraction of tasks in a group that are successful. Coverage is the fraction of tasks in a group for which the agent produces any final answer that can be checked by the automatic evaluator, regardless of whether the answer is ultimately judged correct. As shown in Figure 6(b), we report the task success rate and coverage. Adding Code2MCP tools increases both metrics across all task groups, indicating that a richer MCP tool pool can be effectively exploited by existing planning and retrieval strategies.
@@
Overall, RepoMaster and OpenAgents handle tool discovery and code retrieval on the consumer side, while Code2MCP supplies additional MCP services on the supply side. Under the same front-end selection and planning strategies, integrating Code2MCP enriches the tool pool and yields measurable improvements.
4.5. Case Studies: Protein, Math, and Computational Fluid Dynamics
@@
This subsection highlights three representative repositories from the 50-repository evaluation set: the biomedical protein modeling library ESM, the symbolic mathematics library SymPy, and the CFD framework Foam-Agent (Yue et al., 2025b) built on OpenFOAM, illustrating the kinds of MCP tools Code2MCP generates and how agents use them in practice.
@@
Bridging protein science with AI agents. In protein science, models such as AlphaFold (Jumper et al., 2021) have greatly improved structure prediction, but using them in everyday research still often requires substantial scripting and environment setup. ESM models complement AlphaFold with efficient sequence modeling and zero-shot variant effect prediction, yet many common analyses remain multi-step Python pipelines for computing physicochemical properties, predicting structures, and interpreting mutations. This workflow can be distracting for researchers who primarily focus on wet-lab experiments and want results that are easy to reproduce and explain. Code2MCP converts ESM into an MCP service by exposing tools such as Analyze_sequence, Predict_structure, and Predict_variant_effect. Figure 8(a) illustrates a qualitative example. When a user provides a protein sequence and a mutation site like “Q145G”, the agent can first call Analyze_sequence to produce a baseline characterization, including molecular weight, amino acid composition, and hydrophobicity. It then calls Predict_structure to generate a 3D structure and uses a Predicted Aligned Error (PAE) plot to assess reliability. Finally, Predict_variant_effect helps locate the mutation on the structure and summarize likely impacts, and it can also generate a simple visualization to clarify the mutation’s spatial context.
@@
Enhancing mathematical reliability in AI agents. For mathematical computing, large language models are prone to errors in symbolic derivations and exact calculations, whereas libraries such as SymPy already provide reliable implementations. Code2MCP organizes common capabilities from SymPy into a mathematical MCP service with tools for limits and integration, matrix operations, symbolic simplification, and transforms such as Fourier and Laplace. For example, to find the indefinite integral of $y=\ln(x^{2}+1)$, it calls the corresponding integration tool, and for the volume of revolution of $y=\ln(x)\cdot\sin(x)$ from $x=1$ to $x=2\pi$, it invokes the volume tool to produce the exact result. By grounding sensitive mathematical steps in validated code rather than prompt-based manipulation, the workflow becomes easier to verify and reproduce.
@@
Automating CFD simulation for AI agents. CFD workflows typically involve mesh generation, solver configuration, execution, and visualization, traditionally coordinated by engineers through shell scripts and configuration files—an error-prone and labor-intensive process. Foam-Agent provides a higher-level interface to OpenFOAM, and Code2MCP further turns it into an MCP service with tools such as Generate_mesh, Generate_simulate, and Visualize_velocity, allowing an agent to run the workflow step-by-step in dialogue. As illustrated in Figure 8(b), a user describes the target geometry and physical parameters in natural language, and the agent calls these tools in sequence to construct the mesh, run the simulation, and visualize the velocity field.
5. Conclusion
@@
This paper addresses the key challenge of insufficient tool supply in the MCP ecosystem by introducing Code2MCP, a framework that automatically converts GitHub repositories into functional MCP services. The framework employs a multi-agent workflow for code comprehension, environment reconstruction, and tool abstraction, augmented by a “Run-Review-Fix” self-correcting loop that improves end-to-end reliability across diverse scientific libraries. An evaluation on 50 repositories demonstrates that Code2MCP can robustly expose high-value APIs as MCP tools while also revealing clear failure modes and boundary conditions. Overall, the results highlight an automated pathway for connecting existing codebases to the agent tool ecosystem, while surfacing concrete limitations that delimit the feasibility of fully automated tool supply.
@@ Appendix A Appendix
A.1. Use of Large Language Models
@@
A.2. algorithm
@@
Appendix B Example MCP Tool Implementations Generated by Code2MCP
B.1. Tools Generated from the ESM Repository
MCP Tool for Protein Sequence Analysis @mcp.tool(name="analyze_sequence", description="Analyze protein sequence features (validated, no placeholder fallback).") def analyze_sequence(sequence: str): """ Analyze physicochemical properties of a protein sequence. Returns real computed values; never uses placeholder fallbacks. """ import re import time start = time.time() try: aa_set = set("ACDEFGHIKLMNPQRSTVWY") seq = re.sub(r"[^A-Za-z]", "", sequence or "").upper() seq = "".join([c for c in seq if c in aa_set]) if not seq: return {"success": False, "result": None, "error": "Empty/invalid protein sequence after filtering."} composition = {$a$: seq.count($a$) for $a$ in sorted(aa_set) if seq.count($a$) > 0} try: from Bio.SeqUtils.ProtParam import ProteinAnalysis # type: ignore except Exception as e: return { "success": False, "result": None, "error": ( "Biopython is required for physicochemical analysis but is not installed. " "Install: pip install biopython. " f"Original import error: {str(e)}" ), } pa = ProteinAnalysis(seq) molecular_weight = float(pa.molecular_weight()) isoelectric_point = float(pa.isoelectric_point()) instability_index = float(pa.instability_index()) aromaticity = float(pa.aromaticity()) gravy = float(pa.gravy()) # hydrophobicity helix, turn, sheet = pa.secondary_structure_fraction() sec_struct = {"helix": float(helix), "turn": float(turn), "sheet": float(sheet)} duration_ms = int((time.time() - start) * 1000) result = { "sequence": seq, "length": len(seq), "composition": composition, "molecular_weight": molecular_weight, "isoelectric_point": isoelectric_point, "instability_index": instability_index, "aromaticity": aromaticity, "gravy": gravy, "secondary_structure_fraction": sec_struct, "metadata": {"backend": "biopython_protparam", "duration_ms": duration_ms}, } return {"success": True, "result": result, "error": None} except Exception as e: return {"success": False, "result": None, "error": f"Error during sequence analysis: {str(e)}"}
[…]
B.1.1. Tools Generated from the SymPy Repository
B.1.2. Tools Generated from the Foam-Agent Repository
@@ Appendix C Per-repository Results
Table 3. Repository identifiers, domains, and final outcomes.
@@
  B1   Biomedical   facebookresearch/esm   success   –
@@
  B4   Biomedical   pysam-developers/pysam   success   –
  B5   Biomedical   deepchem/deepchem   success   –
@@
  P2   Psychology   pymc-devs/pymc   success   –
  P3   Psychology   sahahn/BPt   success   –
@@
  P5   Psychology   neuropsychology/NeuroKit   success   –
  M1   Math   sympy/sympy   success   –
@@
  M3   Math   fredrik-johansson/mpmath   success   –
  M4   Math   cvxpy/cvxpy   success   –
  M5   Math   sagemath/sage   success   –
  Ea1   Earth Science   obspy/obspy   success   –
@@
  Ea4   Earth Science   mapbox/rasterio   success   –
  Ea5   Earth Science   geopandas/geopandas   success   –
  A1   Astronomy   astropy/astropy   success   –
@@
  A3   Astronomy   lightkurve/lightkurve   success   –
  A4   Astronomy   astroML/astroML   success   –
  A5   Astronomy   astropy/astroquery   success   –
  C1   Chemistry   rdkit/rdkit   success   –
@@
  C3   Chemistry   bjodah/chempy   success   –
@@
  C5   Chemistry   cclib/cclib   success   –
  Ph1   Physics   csml-rpi/Foam-Agent   success   –
@@
  Ph3   Physics   pydy/pydy   fail   –
  Ph4   Physics   PyAbel/PyAbel   success   –
@@
  S1   Social Science   networkx/networkx   success   –
  S2   Social Science   igraph/python-igraph   success   –
@@
  S5   Social Science   datamade/dedupe   success   –
  L1   Linguistics   nltk/nltk   success   –
  L2   Linguistics   explosion/spaCy   success   –
  L3   Linguistics   stanfordnlp/stanza   success   –
@@
  L5   Linguistics   flairNLP/flair   success   –
  E1   Econometrics   statsmodels/statsmodels   success   –
  E2   Econometrics   alkaline-ml/pmdarima   success   –
@@
  E5   Econometrics   pydata/xarray   success   –
Table 4. Per-repository structural statistics (size, dependencies, and tests).
@@
Table 5. Per-repository environment and MCP conversion statistics.
@@
Appendix D Integration Task Set
This part documents the task set used in the integration experiments with RepoMaster and OpenAgents. The goal of this task set is to evaluate a practical integration question: when Code2MCP supplies MCP services from repositories, can existing agent systems (1) discover the relevant service/tool, (2) invoke it with valid inputs, and (3) complete tool-usage workflows with stable, non-trivial outputs under the same interaction and runtime budgets as the baselines. We intentionally keep the task set focused on integration behavior (selection + invocation + composition) rather than using it as a comprehensive benchmark of general agent capability.
Task sources and construction. We construct a fixed set of tasks by first collecting common task types used in prior agent and tool-usage evaluations (e.g., API-oriented code understanding, single-tool execution, and multi-step tool composition), and then instantiating each type into concrete prompts that are executable and checkable in our MCP setting. We avoid embedding method-specific hints: task statements do not reference Code2MCP, do not assume any particular tool names, and do not provide privileged repository internals beyond what an agent can obtain through its normal repository discovery interfaces. To ensure coverage of both general-purpose and scientific settings, we include a portion of domain-oriented tasks that require using computational libraries via MCP tools, while keeping the task phrasing at the level of user intent rather than narrative case-study descriptions.
What a task looks like. Each task is specified in a uniform template with: (1) a natural-language goal, (2) allowed inputs , (3) constraints relevant to tool usage, and (4) an explicit success condition that is verifiable from tool call traces and returned outputs. The success condition is defined at the tool level: a task is successful only if the agent completes the required tool invocations and the final invocation returns a schema-valid JSON object with three fields (success, result, error), where success is true and result is non-empty/non-null and not a direct echo of the input. This rule is used to rule out superficial “success” responses that do not reflect meaningful tool execution.
Task coverage. The task set spans three interaction patterns that commonly arise in MCP-based workflows. First, tool discovery and selection tasks require the agent to identify which tool corresponds to a stated goal and to infer the minimal valid input format from tool metadata. Second, single-step invocation tasks require a single correct tool call that produces a concrete computed output under the environment constraints. Third, multi-step composition tasks require two or more tool invocations where later inputs depend on earlier outputs, testing whether intermediate results are correctly propagated and whether the agent can recover from minor argument mismatches by adjusting inputs based on returned error messages. We include a small number of “negative” tasks only to probe robustness of invocation behavior; they are evaluated with the same success rule.
  • Low-dependency libraries: at most 10 Python dependencies inferred from requirements.txt or environment.yml, and primarily library-style public APIs.
Evaluation protocol and fairness controls. All methods are evaluated on the same fixed task set with identical task statements and success criteria. We do not modify RepoMaster or OpenAgents planning algorithms for individual tasks, and we do not provide additional hints to methods that use Code2MCP-generated tools. Each run is executed with the same global budgets (maximum interaction steps and wall-clock time) and with the same runtime restrictions used elsewhere in the paper. We log (1) the selected tools, (2) all tool inputs/outputs, (3) per-call durations, and (4) failure reasons (timeouts, schema violations, empty results, or runtime errors). When reporting aggregate results, we compute task success rate as the fraction of tasks that meet the success condition within budget, and we additionally report efficiency metrics consistent with the main paper.
Appendix E Agent Roles and System Prompts
This section outlines the roles of the specialized agents within the Code2MCP framework.
Environment System Prompt - Prefer Conda; use venv only if Conda is unavailable or clearly unsuitable. - Detect dependency sources by priority: environment.yml > requirements.txt > pyproject.toml > setup files; never guess hidden dependencies. - Pin versions when explicit; otherwise install the minimal viable set. Prefer CPU wheels unless GPU is explicitly required. - Normalize cross-platform behavior; avoid absolute paths; use relative POSIX-like paths; ensure UTF-8 locale. - Smoke test: print Python version and platform; import fastmcp; attempt to import the project’s top-level package or a primary CLI; exit code 0 indicates pass. - On failure, capture exact command, exit code, last 80 lines of error logs, and timing; propose exactly one minimal remedy (e.g., switch to venv, install a single missing package, try one version pin, extend timeout once). - Apply at most one fallback; never change repository code; do not write outside the workspace; do not weaken security (e.g., no SSL bypass). - Cache wheels where possible; avoid global pollution; record reproducible commands and resolved versions. - Default to offline validation; if network is strictly required, justify briefly and bound the scope. - Emit a compact environment report (platform, Python, manager, explicit deps, resolved pins, pass/fail).
Review & Auto-Fix System Prompt - Triage failures: import/env, type/contract, path/I-O, dependency/version, timeout/perf, platform. - Choose minimal direct fix vs. regeneration; provide a one-line rationale and confidence (low/med/high). Prefer adapter-boundary mitigations (lazy import, existence checks, parameter coercion). - Apply strict complete-file replacement for the single target file; return pure code only; do not alter unrelated sections. - Preserve interface contracts and standardized error shapes; add narrow guards instead of broad catches. - Enforce cross-platform path handling and deterministic behavior; do not introduce external network calls or new side effects. - Optionally add a tiny internal sanity check if it prevents recurrence without bloat. - Bound attempts (<= B). If still failing, emit a compact escalation note: failing step, last command, error logs tail, and the next best single remediation.
E.1. GPT-4 Template Baseline Prompt